Air at a pressure of 100 kPa has a volume of 0.32 m3. The air is compressed in a reversible adiabatic manner until the temperature is 190 oC. The reversible work is -63 kJ/kg. Determine the initial temperature.

t =


1
Expert's answer
2021-10-08T09:24:45-0400

For air "k=1.4,"

"A=\frac{p_1V_1}{k-1}(1-(\frac{V_1}{V_2})^{k-1}),\implies" "V_2=\frac{V_1}{1-(\frac{(k-1)A}{p_1V_1})^\frac 1{k-1}},"

"T_1V_1^{k-1}=T_2V_2^{k-1},\implies"

"T_1=\frac{T_2}{1-(\frac{(k-1)A}{p_1V_1})^\frac 1{k-1}}=208~K," or

"t_1=-65\u00b0C."